OpenStudy (anonymous):

Let f(x) = -20x2 + 14x + 12 and g(x) = 5x - 6. Find F/G and state its domain. a. –4x – 2; all real numbers except x =6/5 b. 5x - 6; all real numbers c. –4x – 2; all real numbers d.5x - 6; all real numbers except x =6/5

OpenStudy (neer2890):

f(x)=-20x^2+14x+12 =-20x^2-10x+24x+12 =-10x(2x+1)+12(2x+1) =(-10x+12)(2x+1) g(x)=5x-6\[\frac{ f(x) }{ g(x) }=\frac{ (-10x+12)(2x+1) }{ (5x-6) }\]

OpenStudy (neer2890):

\[\frac{ f(x) }{ g(x) }=\frac{ -2(5x-6)(2x+1) }{ (5x-6) }\] \[=-2(2x+1)\] =-4x-2

OpenStudy (neer2890):

to find domain of f(x)/g(x), if (5x-6)=0 then it will give divide by 0 error. so,\[(5x-6) \neq 0\] \[x \neq \frac{ 6 }{ 5 }\]

OpenStudy (neer2890):

so domain is all real numbers except 6/5 and f(x)/g(x)=-4x-2

OpenStudy (anonymous):

thanks man sorry for late response I left school before having the chance to respond could you possibly answer some other math questions i may have?

OpenStudy (neer2890):

you're welcome.. and yes definitely,i'm here to help...:)

OpenStudy (anonymous):

well look I have like 5 more test to do which are like 20 questions each (algebra 2) on GradPoint.com if you can do all those tests I would happily pay you 30$ or just name the price......on paypal could you possibly do this for me?